OpenStudy (anonymous):

Algebra Help Write each function in intercept form. f(X)=-3x^2+24x-45

OpenStudy (anonymous):

Here does this help? It explains better than I could. http://blogformathematics.blogspot.com/2011/03/intercept-form-of-quadratic-equation.html

OpenStudy (anonymous):

It does. Thank you.

OpenStudy (anonymous):

Well actually that's not true. If we just factor the quadratics, we can find the intercept form duh

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

So f(x) = 2x^2-8x+6 y = 2(x^2-4x+3) y = 2(x-3)(x-1)

OpenStudy (anonymous):

Where y = 2(x-3)(x-1) is the intercept form of f(x) = 2x^2-8x+6

OpenStudy (anonymous):

Oh. Thank you. I understand it now.

OpenStudy (anonymous):

I still need help with the second one.

OpenStudy (anonymous):

I took first one out of the question.

OpenStudy (anonymous):

-3x^2+24x-45=-3(x^2+8x-15) I think you can continue alone from here

OpenStudy (cwrw238):

take out -3 and factor

OpenStudy (anonymous):

Would it be 5 and 3?

OpenStudy (anonymous):

Im not sure how to factor it.

OpenStudy (anonymous):

use quadratic formula

OpenStudy (cwrw238):

is that a typo @myko shouldn't it be + 15?

OpenStudy (cwrw238):

and -8

OpenStudy (anonymous):

then it would be -5 and -3

OpenStudy (anonymous):

or notice that: \((x-a)(x-b)=x^2-(a+b)x+ab\) so in your case: \(-(a+b)=8\) and \(ab=-15\) solve this for a and b

OpenStudy (anonymous):

So the answer would be f(x)=-3(x-5)(x-3) Right?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Okay. Thank y'all for the help.

OpenStudy (cwrw238):

yw